cos 35 cos 85 cos 155=?
 
 

Grade:12th pass

1 Answers

Lab Bhattacharjee
8 years ago
cos(155)=cos(180-25)=-cos(25)
 
So, cos35 cos85 cos155 =-cos25 cos(60-25)cos(60+25)
 
Now prove 4cosx cos(60-x) cos(60+x)=cos3x
 
cos75=cos(45+30)=...